OpenStudy (anonymous):

Jessica has taken 6 English tests and scored: 87, 92, 85, 85, 89, and 98. What does she need to make on her next test to have an A, or an average of 90, in the class? A. 4 B. 54 C. 94 D. 100

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

94 right?

OpenStudy (bruno102):

Yes correct!

OpenStudy (shamil98):

(87 + 92 +85 +85 +89 +98 + x )/ 7 = 90

OpenStudy (shamil98):

hold on -.-

OpenStudy (shamil98):

536 + x / 7 = 90 536 + x = 630 x = 94 so yes you are correct.